Hi Tutors:
Word prolems are a work in progress for me. Please show me how to work this one:
A business club split a $30,000 investment among three bond funds. The first one grew at 2%, the second at 3% and the third at 6%. The total earnings were $1,110. The earnings from the second fund were $150 more than those of the first. How much was invested in each fund?

Answer by longjonsilver(2297)

x+y+z = 30000 --eqn1

0.02x + 0.03y + 0.06z = 1110 --eqn2

0.03y = 0.02x + 150 --eqn3

OK... multiply eqn 1 by 0.06 then subtract eqn2, leaving 0.04x + 0.03y = 690 --eqn4

Now rearragne eqn3 around and then add it to eqn4, giving x = 9000.

Then find y and z as per the other Q.

jon.